For each of the following, compute the present value:

Present Value

Years

Interest Rate

Future Value

15

7%

$19,415

8

11

47,382

13

10

312,176

25

13

629,381
Answer #1

rate positively ..

Computation of present value
Present value
1 PV = 19,415/((1.07)^15)              7,036.89
2 PV = 47,382/((1.11)^8)            20,560.31
3 PV = 312176/((1.1)^13)            90,426.27
4 PV = 629,381/((1.13)^25)            29,645.07